OpenStudy (anonymous):

The measure of an angle is 51°. What is the measure of its supplement? A. 39° B. 49° C. 129° D. 139°

hartnn (hartnn):

Supplement of angle 'x' is just 180-x

jimthompson5910 (jim_thompson5910):

two supplementary angles always add to 180 degrees x+y = 180 51+y = 180 y = ??

OpenStudy (anonymous):

C

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

Thanks :) 2 more?

hartnn (hartnn):

sure

OpenStudy (anonymous):

The lines intersect at point K. http://static.k12.com/calms_media/media/1421000_1421500/1421001/1/461d4a8736bac521aee4fecbb679517fbeac36b2/MS_Alg_130924_281813.jpg What is the value of x? A. 23 B. 30 C. 46 D. 62

hartnn (hartnn):

vertically Opposite angles are congruent.

hartnn (hartnn):

5x-53 = 3x-7 can you find 'x' from here ?

OpenStudy (anonymous):

um... B?

OpenStudy (anonymous):

... um.... ok?

hartnn (hartnn):

lol, did you even calculate? or it was just your guess ? because B is correct :P

OpenStudy (anonymous):

that was a guess XD but I understand that I needed to multiply and make it match :D

hartnn (hartnn):

oh wait, B is not correct

OpenStudy (anonymous):

I didn't get B

hartnn (hartnn):

5x-53 = 3x-7 2x = 53-7 = 46 x=23

OpenStudy (anonymous):

oh... oops sorry

OpenStudy (anonymous):

That's what I got too :P

hartnn (hartnn):

yeah, sometimes when you solve a problem mentally, such silly mistakes happen :P

OpenStudy (anonymous):

I solved and got A mentally, but mistakes happen.. specially when they really matter xD

OpenStudy (anonymous):

but thanks for noticing that XD but how would I do this one?

OpenStudy (anonymous):

Ok first you must find x, and by looking at the picture you can find 8x+9x+5=90

hartnn (hartnn):

angle BHR and angle NHD are vertically opposite angles

OpenStudy (anonymous):

wait I got this 1 sec :D

OpenStudy (anonymous):

nvm continue explaining

hartnn (hartnn):

i'll let venny take this one :)

OpenStudy (anonymous):

And what @hartnn said, angle BHR and NHD are verically opposite so the angle we need to find is equal to 8x

OpenStudy (anonymous):

Have you solved for x yet? or do you need help with that

OpenStudy (anonymous):

I need help sorry

OpenStudy (anonymous):

Its okay! 8x+9x+5=90 The first step is to combine like terms, 17x+5=90 next you need to subtract 5 so 17x=85 Divide both sides by 17

OpenStudy (anonymous):

what do you get for x?

OpenStudy (anonymous):

1x=5?

OpenStudy (anonymous):

Yes! and since angle NHD=angleBHR, we know that the angle we are trying to find is equal to 8x, So to find that we multiply 5(8)

OpenStudy (anonymous):

40 so its A?

OpenStudy (anonymous):

Yep! I got A too

OpenStudy (anonymous):

XD I FEEL GREAT! Thank you soo much :)

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

thanks to @hartnn too :)

OpenStudy (anonymous):

indeed for showing me how to solve the 1st one, thanks everyone! :) :D XD

hartnn (hartnn):

welcome ^_^ happy to help :)

OpenStudy (anonymous):

th edog of os is @hartnn

OpenStudy (anonymous):

can I hit you guys up later if I need the help?

OpenStudy (anonymous):

th edog of os is @hartnn

OpenStudy (anonymous):

th edog of os is @hartnn

OpenStudy (anonymous):

sure if i'm not asleep

OpenStudy (anonymous):

th edog of os is @hartnn

OpenStudy (anonymous):

ok XD

OpenStudy (anonymous):

th edog of os is @hartnn

OpenStudy (anonymous):

th edog of os is @hartnn

hartnn (hartnn):

wow, i am obliged , thanks foooo

OpenStudy (anonymous):

th edog of os is @hartnn

OpenStudy (anonymous):

th edog of os is @hartnn

OpenStudy (anonymous):

th edog of os is @hartnn

OpenStudy (anonymous):

I would give you a badge for that @hartnn , but I already did

OpenStudy (anonymous):

So much spammers on here recently :(

hartnn (hartnn):

don't get sad lol, just ignore them :)